OpenStudy (anonymous):

Suppose Kristin buys 10 tickets out of 150 which will award 5 prizes on the basis of a random drawing. What are the chances that Kristin will win: a.) no prize? b.) exactly 3 prizes? c.) all the 5 prizes? d.) at least one prize?

OpenStudy (mathmate):

Hints: This is an appropriate application of the binomial probability distribution. let n=number of prizes (5) p=probability of winning (10/150) r=number of prizes won P(r)=probability of winning r prizes then \(P(r)=nCr~ p^r (1-p)^{n-r}\) and \(nCr = \dfrac{n!}{r!(n-r)!}\) You can then proceed to calculate the P(r) for the following cases: a) r=0 b) r=3 c) r=5 d) 1-P(0)

OpenStudy (anonymous):

Ohhhh. I don't understand binomial probability distribution. Can you help me a little further? :)

OpenStudy (mathmate):

@killua_vongoladecimo Do you mean you are Not supposed to use binomial, or you need help with binomial?

OpenStudy (anonymous):

I need help with binomial. Hmm So let me give it a try for A.) \[P \left( r \right) = 5C0 (\frac{ 1 }{ 15 })^{0} (\frac{ 14 }{ 15 })^{5}\] Is this right?

OpenStudy (anonymous):

So here are the values I got: a.) 0.708245596 b.) 0.002581069959 c.) 0.000001316872428 d.) 0.291754403

OpenStudy (mathmate):

All correct and well done. :)